solve by substitution y=2x+1 the second part is 3x+y=11
Found 2 solutions by ikleyn, fractalier:
Answer by ikleyn(52824)
Answer by fractalier(6550)

Let us solve this one for you...
y=2x+1
3x+y=11
Plugging the first one into the second gives you
3x + (2x+1) = 11
3x + 2x + 1 = 11
5x + 1 = 11
Now sbtract one and divide by five and get
5x = 10
x = 2 so that
y = 2(2) + 1 = 4 + 1 = 5
(2, 5)